Lean Accounting

Modern Lighting Inc. manufactures lighting fixtures, using lean manufacturing methods. Style Omega has a materials cost per unit of $34. The budgeted conversion cost for the year is $172,800 for 3,200 production hours. A unit of Style Omega requires 12 minutes of cell production time. The following transactions took place during June:

  1. Materials were acquired to assemble 770 Style Omega units for June.
  2. Conversion costs were applied to 770 Style Omega units of production.
  3. 720 units of Style Omega were completed in June.
  4. 670 units of Style Omega were sold in June for $90 per unit.

a. Determine the budgeted cell conversion cost per hour. If required, round to the nearest cent.
$ per hour

b. Determine the budgeted cell conversion cost per unit. If required, round to the nearest cent.
$ per unit

Feedback

a. Budgeted conversion cost ÷ Planned hours of production = Cell Conversion cost per hour

b. (Cell process time ÷ 60 minutes) x Conversion rate from Req. (a) = Cell Conversion cost per unit

c. Journalize the summary transactions (1)–(4) for June. If required, round your answers to one decimal place. If an amount box does not require an entry, leave it blank.

1. Raw and In Process Inventory
Accounts Payable
2. Raw and In Process Inventory
Conversion Costs
3. Finished Goods Inventory
Raw and In Process Inventory
4. Sale Accounts Receivable
Sales
4. Cost Cost of Goods Sold
Finished Goods Inventory
Answer #1
a
Budgeted cell conversion cost per hour 54.00 per hour =172800/3200
b
Budgeted cell conversion cost per unit 10.80 per unit =54*(12/60)
c
Raw and in process inventory 26180 =770*34
     Accounts Payable 26180
Raw and in process inventory 8316 =770*10.80
       Conversion costs 8316
Finished goods inventory 32256 =720*(34+10.80)
       Raw and in process inventory 32256
Accounts Receivable 60300 =670*90
       Sales 60300
Cost of goods sold 30016 =670*(34+10.80)
        Finished goods inventory 30016